# Enhancing Data Privacy in Large Language Models through Private Association Editing

### v1

@@ 4.2 Results and Discussion
Since LLMs tend to leak training data, we aim to quantify the amount of private information that can be retrieved from the pre-trained GPT-J. Unfortunately, GPT-J makes no exception to the trend noticed by Huang et al. (2022) for the GPT-Neo models. In fact, also this model tends to generate PII.
@@
Moreover, as the attacker gets more information, the accuracy of the attacks gets higher: the accuracy of the attacks strongly depends on the length of the prompt. In fact, the lower accuracy – the number of correctly leaked email addresses over the total email adresses generated– that can be registered in Memorization Attacks is $12\%$ : the model in that case is fed with a context prompt that is $50$ tokens long. However, when the context prompt given to the model is composed of $200$ tokens, the accuracy of the attack peaks at $18.2\%$ with greedy decoding and $18\%$ using beam search decoding.
The accuracy of the Association Attacks is much more modest. The results of those attacks against GPT-J model exhibit similar patterns to the one observed by Huang et al. (2022) against the GPT-Neo models. The larger number of email addresses leaked by those kind of attacks is $68$, a modest number compared to the accuracy obtained in the Memorization Attacks. However, in an adversarial scenario even low accuracy may cause harm to people. Hence, in the next Section, we will demonstrate the efficacy of PAE against both types of attacks.
@@
Post-edit results show, in fact, a significant reduction in the effectiveness of Association attacks. This reduction is particularly notable in scenarios where the number of leaked emails drops close to zero. For example, under the Implicit strategy, zero-shot b result in 0 leaked emails and 0 Attack accuracy, indicating a complete mitigation of the attack. However, some prompts still cause leakage; for instance, in prompt zero-shot b the edit reduces the number of email addresses leaked significantly but not completely (from 68 pre-edit to 48 post-edit in the Implicit). Crucially, while not perfect, the PAE edits – both Implicit and Explicit – always cause an increase in privacy protection, since reduce the number of email correctly leaked by Association Attacks.
@@
However, it is crucial to consider the originally leaked emails when interpreting post-edit results. While a reduction to near-zero leakage is impressive, the impact is more pronounced when starting from a higher number of pre-edit leaks. For this reason, we focus on the discussion around the Memorization Attacks, that cause a larger number of private email addresses to be generated.
@@
PAE is an effective solution against Memorization Attacks. In particular, the accuracy of the attacks steady decreases in each configuration. The average drop in accuracy after an Implicit edit is $5\%$ and $4.5\%$ after Explicit edit: this means that PAE is able to modify model parameter so that, on average, the $32\%$ of the previously predicted email addresses are no more verbatim generated by the model using Implicit defense strategy, $29\%$ with the explicit one. Against attacks with context prompt of 50 tokens, PAE effectiveness peaks, with $42.5\%$ of the email addresses anonymized. As expected, more informed context prompts are more challenging: however, also with context prompts of 200 tokens, PAE make the accuracy attack drop from $0.18$ to $0.12$ in Greedy decoding and to $0.138$ using Beam Search in the Implicit edit, and from $0.18$ to $0.134$ in greedy decoding and to $0.141$ using Beam Search in the Explicit edit. In general, studying the effect of the decoding algorithm on the attacks accuracy we can state that this factor does not influence much the results: under Memorization Attacks only a slight difference in term of accuracy can be registered From this analysis can conclude that PAE can help in protecting privacy.
Finally, it is possible to notice that there is a consistent difference between Implicit and Explicit post-edit results. Explicit edits generally result in a slightly higher number of leaked emails and attack accuracy compared to Implicit edits, expecially under Memorization Attacks. For example, in the case of a context prompt of 200, the Explicit edit cause a larger number of emails to be correctly generated (396 in greedy deconding, 405 in beam search decoding) than the corresponding Implicit edit (368 email addresses leaked in greedy decoding, 394 in beam search decoding)
@@
In summary, post-edit measures, particularly implicit edits, demonstrate a strong capability to safeguard email data from various attack strategies, significantly lowering both the number of leaked emails and the attack accuracy across different configurations. For our experiments, we adopted the "one model, n edits" philosophy. This approach is based on the practical scenario where an LLM producing private data needs to be edited for a large number of potentially threatened individuals: with PAE, the model owner can perform a single edit to the model parameters to reduce privacy risks. Our investigation on the emails that the model generates when subjected to Memorization and Association attacks confirm that the Memorization Attacks are able to recover a larger number of private information also after the editing. However, the evaluation confirms the effectiveness of our model editing techniques in preventing the disclosure of private data since also informed attacks –like the Memorization Attacks – are less effective on edited models. This setup is particularly challenging because it requires analyzing the impact of n modifications at the same time. In the next Section, we also demonstrate that the original language model is not negatively influenced by PAE.
@@
$E_{1}$-$O$ $0.808(\pm 0.198)$
$E_{2}$-$O$ $0.793(\pm 0.199)$
BLEU $E_{1}$-$E_{2}$ $0.790(\pm 0.203)$
$E_{1}$-$O$ $0.841(\pm 0.173)$
$E_{2}$-$O$ $0.826(\pm 0.172)$
Automatic Evaluation METEOR $E_{1}$-$E_{2}$ $0.824(\pm 0.183)$
Accuracy $0.35(\pm 0.07)$
Manual Revision Fleiss’ $K$ $0.002$
Table 3: PAE preserves language model performances. In the Manual revision of the generation, annotators are not able to detect which model generated each paragraph. Moreover, on a larger scale of examples, the Automatic Evaluation reveals that the generation of post-edit LLMs $E_{1}$ and $E_{2}$ are both similar to one another ($E_{1}$-$E_{2}$) and with respect to the pre-edit LLM $O$ ($E_{1}$-$O$ and $E_{2}$-$O$)
@@
PAE preserves privacy of people while not affecting the Language Modeling performances. Results of the evaluation can be found in the Table 3. The results of the automatic assessment procedure can quantitatively give us insight that models generations are, in fact, similar. Both according to BLEU metric and to METEOR, the systems generate (in greedy decoding) very similar paragraph when prompted with the same tokens. In particular, the post-edited models $E_{1}$ and $E_{2}$ – post-edited with implicit and explicit PAE – are similar to the original, pre-edited model $O$ and are also similar with respet to each other. Finally, the manual assessment procedure suggest that the models are indistinguishable from one another. In fact, the annotators asked to detect wich model is responsible for a generation among $E_{1}$, $E_{2}$, and $O$ can only randomly guess, with an average accuracy on this classification task ($0.35(\pm 0.07)$) close to random choice. Also the very low agreement suggest that tge three systems are indistinguishable.
@@
This evaluation procedure can attest that the EAP is applicable because it not only helps to preserve user privacy, but also leaves the capabilities of the systems language model intact.
@@
Figure 2: Memorization Attack against sequentially post-edited models. The smaller the batch size $k$, the larger the number of sequential updates are necessary to edit all the private emails addressess leaked by the original model. After the Sequential edits, the stronger Memorization Attack ($|p_{M}|=200$) achieve similar performances at all the configurations.
PAE is applicable with sequential editing
@@
Finally, in Table is it possible to notice that the sequential update is definetely applicable with PAE. In this experiments, we perform sequential edit of the GPT-J model, varying the number of email addresses anonymized per edit, varying from $5$ to $300$. We indicate the number of anonymized emails per edit as batch size $k$: with $k<N$ we mimic the real world scenario of updating a model each time a privacy leak is detected. To understand whether sequential editing has a negative impact on effectiveness of the edit, we evaluate the effectiveness of PAE for each of the batch sizes in the Memorization Attack with the more effective of the prompts ($|p_{M}|=200$). The results in Table 4.2 refers to a model post-edited with "implicit" PAE. As can be observed in Table 4.2, the accuracy of the edit is rather stable and similar to the results obtained in the batch editing scenario. Those results confirm the applicability of PAE also in sequential editing.
@@ 5 Conclusion
In this paper, we address the critical issue of private data leakage in Large Language Models (LLMs) due to their tendency to memorize training data. We propose Private Association Editing (PAE), a novel defense mechanism that effectively removes Personally Identifiable Information (PII) from LLMs without requiring retraining.
Our methodology involves a four-step procedure: detecting memorized PII, applying PAE cards, verifying resilience to targeted data extraction (TDE) attacks, and ensuring consistency in the post-edit LLMs. The PAE method stands out for its versatility and efficiency, allowing for small batch modifications and significantly enhancing the privacy of LLMs.
Our experiments demonstrate that the PAE approach is both effective and efficient in mitigating the risk of private data leakage. We believe PAE will be a valuable tool in the ongoing effort to protect data privacy in LLMs and encourage its adoption to prevent potential privacy violations as these models continue to be deployed in real-world applications.

### v2

@@ 4.2 Results and Discussion
Since LLMs tend to leak training data, we aim to quantify the amount of private information that can be retrieved from the pre-trained GPT-J. Unfortunately, GPT-J makes no exception to the trend noticed by Huang et al. (2022) for the GPT-Neo models. In fact, this model also tends to generate PII.
@@
Moreover, as the attacker gets more information, the accuracy of the attacks gets higher: the accuracy of the attacks strongly depends on the length of the prompt. In fact, the lower accuracy – the number of correctly leaked email addresses over the total email addresses generated– that can be registered in Memorization Attacks is $12\%$ : the model in that case is fed with a context prompt that is $50$ tokens long. However, when the context prompt given to the model is composed of $200$ tokens, the accuracy of the attack peaks at $18.2\%$ with greedy decoding and $18\%$ using beam search decoding.
The accuracy of the Association Attacks is much more modest. The results of those attacks against GPT-J model exhibit similar patterns to the one observed by Huang et al. (2022) against the GPT-Neo models. The larger number of email addresses leaked by those kinds of attacks is $68$, a modest number compared to the accuracy obtained in the Memorization Attacks. However, in an adversarial scenario, even low accuracy may cause harm to people. Hence, in the next Section, we will demonstrate the efficacy of PAE against both types of attacks.
Wikipedia Books3 Pile-CC
BLEU METEOR BLEU METEOR BLEU METEOR
$O$ $E_{1}$ 0.715(±$(\pm 0.266)$ 0.754(±$(\pm 0.237)$ 0.66(±$(\pm 0.295)$ 0.71(±$(\pm 0.259)$ 0.599(±$(\pm 0.27)$ 0.656(±$(\pm 0.249)$
$E_{2}$ 0.683(±$(\pm 0.277)$ 0.733(±$(\pm 0.247)$ 0.589(±$(\pm 0.276)$ 0.646(±$(\pm 0.247)$ 0.533(±$(\pm 0.243)$ 0.601(±$(\pm 0.222)$
$E_{1}$ $E_{2}$ 0.662(±$(\pm 0.272)$ 0.713(±$(\pm 0.239)$ 0.63(±$(\pm 0.285)$ 0.679(±$(\pm 0.253)$ 0.591(±$(\pm 0.267)$ 0.65(±$(\pm 0.25)$
Table 3: PAE preserves language model performances:generations from the original, pre-edit model ($O$) and post-edit ($E_{1}$ and $E_{2}$) are similar from one another.
@@
Post-edit results show, in fact, a significant reduction in the effectiveness of Association attacks. This reduction is particularly notable in scenarios where the number of leaked emails drops close to zero. Crucially, while not perfect, the PAE edits – both Implicit and Explicit – always cause an increase in privacy protection since they reduce the number of emails correctly leaked by Association Attacks. However, it is crucial to consider the originally leaked emails when interpreting post-edit results. While a reduction to near-zero leakage is impressive, the impact is more pronounced when starting from a higher number of pre-edit leaks. For this reason, we focus on the discussion around the Memorization Attacks that cause a larger number of private email addresses to be generated.
@@
PAE is an effective solution against Memorization Attacks. In particular, the accuracy of the attacks steadily decreases in each configuration. The average drop in accuracy after an Implicit edit is $5\%$ and $4.5\%$ after an Explicit edit: this means that PAE is able to modify model parameters so that, on average, the $32\%$ of the previously predicted email addresses are no longer verbatim generated by the model using Implicit defense strategy, $29\%$ with the explicit one. Against attacks with context prompt of 50 tokens, PAE effectiveness peaks, with $42.5\%$ of the email addresses anonymized. As expected, more informed context prompts are more challenging; however, also with context prompts of 200 tokens, PAE makes the accuracy attack drop from $0.18$ to $0.12$ in Greedy decoding and to $0.138$ using Beam Search in the Implicit edit, and from $0.18$ to $0.134$ in greedy decoding and to $0.141$ using Beam Search in the Explicit edit. In general, studying the effect of the decoding algorithm on the accuracy of the attacks we can state that this factor does not influence much the results: under Memorization Attacks, only a slight difference in terms of accuracy can be registered This analysis show that PAE can help in protecting privacy. Finally, it is possible to notice that there is a consistent difference between Implicit and Explicit post-edit results. Explicit edits generally result in a slightly higher number of leaked emails and attack accuracy compared to Implicit edits, especially under Memorization Attacks. For example, in the case of a context prompt of 200, the Explicit edit causes a larger number of emails to be correctly generated (396 in greedy decoding, 405 in beam search decoding) than the corresponding Implicit edit (368 email addresses leaked in greedy decoding, 394 in beam search decoding)
@@
In summary, post-edit measures, demonstrate a strong capability to safeguard email data from various attack strategies, significantly lowering both the number of leaked emails and the attack accuracy across different configurations. For our experiments, we adopted the “one model, $k$ edits” philosophy. This approach is based on the practical scenario where an LLM producing private data needs to be edited for a large number of potentially threatened individuals: with PAE, the model owner can perform a single edit to the model parameters to reduce privacy risks. While attacks are still able to recover some private information also after the editing, the evaluation confirms the effectiveness of our model editing techniques in preventing the disclosure of private data since also informed attacks –like the Memorization Attacks – are less effective on edited models.
@@
PAE preserves the privacy of people while not affecting the Language Modeling performances. The results of the automatic assessment procedure can quantitatively give us insight that model generations are, in fact, similar. Results can be found in the Table 3. Both according to the BLEU metric and to METEOR, the systems generate (in greedy decoding) very similar paragraphs when prompted with the same tokens. In particular, the post-edit models $E_{1}$ and $E_{2}$ – post-edit with implicit and explicit PAE – are similar to the original, pre-edited model $O$. In particular, the implicit edit $E_{1}$ is slightly more similar to the pre-edited model across all the domains tested.
Finally, the manual assessment procedure suggests that the models are indistinguishable from one another. In fact, the annotators asked to detect which model is responsible for a generation among $E_{1}$, $E_{2}$, and $O$ can only randomly guess, with an average accuracy on this classification task ($0.35(\pm 0.07)$) close to random choice. Also, the very low agreement (Fleiss’ K of $0.002$) suggests that the three systems are indistinguishable.
@@
This evaluation procedure can attest that the EAP is applicable because it not only helps to preserve user privacy but also leaves the capabilities of the systems language model intact.
PAE “one model, $k$ edits” is flexible
@@
Figure 2: Memorization Attack against sequentially post-edit models. The smaller the batch size $k$, the larger the number of sequential updates necessary to edit all the private email addresses leaked by the original model. After the Sequential edits, the stronger Memorization Attack ($|p_{M}|=200$) achieves similar performances at all the configurations.
@@
Finally, in Figure 2 is it possible to notice that the “one model $k$ edits approach” is flexible and can be applied with different $k$, successfully mixing batch and sequential editing to preserve users’ privacy. In these experiments, we perform sequential edits of the GPT-J model, varying the number of email addresses anonymized per edit, varying $k$ from $5$ to $300$. We indicate the number of anonymized emails per edit as batch size $k$: with $k\ll N$, we mimic the real-world scenario of updating a model each time a privacy leak is detected. To understand whether sequential editing has a negative impact on the effectiveness of the edit, we evaluate the effectiveness of PAE for each of the batch sizes in the Memorization Attack with the more effective of the prompts ($|p_{M}|=200$). The results in Table LABEL:tab:seqedit refer to a model post-edit with "implicit" PAE. As can be observed in Figure 2, the accuracy of the edit is rather stable and similar to the results obtained in the batch editing scenario. Also, the underlying language model is not negatively affected by the different $k$, as reported in Appendix 6.2. Those results also confirm the applicability of PAE in sequential editing, demonstrating the validity of the “one model, $k$ edits” approach.
@@ 5 Conclusion
In this paper, we address the critical issue of private data leakage in Large Language Models (LLMs) due to their tendency to memorize training data. We propose Private Association Editing (PAE), a novel defense mechanism that effectively removes Personally Identifiable Information (PII) from LLMs without requiring retraining. Our methodology involves a four-step procedure: detecting memorized PII, applying PAE cards, verifying resilience to targeted data extraction (TDE) attacks, and ensuring consistency in the post-edit LLMs. The PAE method stands out for its versatility and efficiency, allowing for small batch modifications and significantly enhancing the privacy of LLMs.
Our experiments demonstrate that the PAE approach is both effective and efficient in mitigating the risk of private data leakage. The results across different configurations suggest its applicability also to other structured PII (i.e. phone numbers and credit cards). We believe PAE will be a valuable tool in the ongoing effort to protect data privacy in LLMs and encourage its adoption to prevent potential privacy violations.